# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v2

@@ 7 Trust and Explainability in Agentic AI
Agentic AI systems are highly autonomous agents capable of making decisions and taking actions without continuous human oversight. These systems also pose unique challenges and opportunities for human trust. Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms. We show the relationship between trust mechanisms, explainability, and evaluation facets in Figure 7 and discuss next.
@@ 7.1 Building Trust in Autonomous Agents
Transparency and Decision Traceability. Transparency is one of the core enablers of trust. For users to understand and evaluate agentic decisions, the reasoning chains, decision states, and action triggers of agents must be made observable. Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent [139] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPT[140] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
@@
Behavioral Consistency and Bounded Autonomy. Trust requires predictability. Agents should follow defined roles, output in expected formats, and remain within delegated authority. For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow [141], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
Social Trust Cues and Language Behaviors. Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [152, 16, 153]. For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries [142, 143]. Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [154]. Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [144, 145]. This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability. These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
@@ 7.2 Explainability Techniques in AMAS
Despite these advancements, achieving robust explainability in Agentic AI systems remains an open research problem. Many techniques focus on isolated predictions or modules and fail to capture system-level dynamics. Future work should prioritize longitudinal interpretability across agent interactions, causal reasoning pipelines, and interactive querying interfaces that support transparency in real time. A comparative summary of current techniques is shown in Table 6.
7.3 Evaluation Metrics for Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI system performance and its real-world implications. Below, we discuss these metrics and summarize in Table 7.
Trustworthiness. This criteria evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [164]. Some approaches combine such factors into an overall trust score. For example, one model defines a trustworthiness score $T$ as
@@ 7.3 Evaluation Metrics for Agentic AI Systems
Coordination. In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
@@
8 Security and Privacy in LLM-based Agentic Multi-Agent Architectures
@@ 8.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [191]. Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [192]. Among the foundational techniques employed are encryption [193, 194], access control [195, 171], adversarial defense [196, 197], and runtime monitoring[198, 199] each adapted to the unique demands of decentralized multi-agent environments.
@@
Encryption. plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [194, 61, 200]. Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption [201, 200], and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
@@
Access control. becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen[18] and CrewAI [81] where agents take on specialized roles (e.g., summarizer, planner, coder) [202], enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [18, 198, 203]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [204, 205] and Attribute-Based Access Control (ABAC) [206, 205] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels [5].
@@
Adversarial robustness. is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results [207, 208]. Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others [208, 192, 115, 55, 116, 209]. Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities [210, 115]. Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
@@
Runtime monitoring. systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity [211]. Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [212, 213, 125]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
As Agentic AI systems continue to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount. Future approaches must include zero-trust frameworks, secure multi-party computation, and formal verification of inter-agent protocols to ensure safe and resilient operation across decentralized autonomous agent collectives.
@@ 8.2 Privacy-Preserving Techniques
The decentralized and interactive nature of LLM-based Agentic AI systems introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory. To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
Differential privacy (DP). DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [214, 215]. DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
@@
Privacy-by-design principles are becoming central to the engineering of next-generation agentic systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
8.3 Compliance and Governance
As Agentic AI systems grow in capability and autonomy, ensuring regulatory compliance and instituting robust governance mechanisms become imperative. Unlike traditional AI agents, Agentic AI systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications. Effective governance in this context spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
@@ 8.3 Compliance and Governance
Regulatory standards. Standards for regulation provide the baseline requirements that all AI systems, including agentic architectures, must meet. Frameworks such as the NIST AI Risk Management Framework (AI RMF) and the EU AI Act define principles for trustworthy AI including transparency, accountability, and fairness. These standards are especially relevant for LLM-based AMAS that interact in high-stakes domains like healthcare, finance, defense, or transportation. For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring of such systems.
Auditability. This is critical in ensuring transparency and facilitating post-hoc accountability. Each decision, plan, or interaction within an agentic system should be logged with timestamps, context, agent role, and justification. Techniques such as decision provenance and action traceability enable this, allowing regulators or internal auditors to reconstruct how decisions were reached. For instance, in systems like AutoGen or MetaGPT, where agents assume specialized roles (e.g., researcher, coder, reviewer), audit trails can capture role-specific actions and flag inconsistencies, bias amplification, or security violations. Blockchain-based audit logs are also being explored to ensure immutability and verifiability of multi-agent interactions.
Policy enforcement. Policy governs what Agentic AI system can and cannot do, and under which conditions. These policies must be codified into the orchestration layers or meta-agent governance modules that manage agent interactions. Examples include enforcing memory expiration policies to avoid data retention violations or restricting access to external tools based on role and authentication level. Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent sub-components. Furthermore, real-time monitoring systems can halt or flag agent activity that deviates from pre-specified ethical constraints or operational bounds, using formal verification tools like TLA+ or symbolic execution engines.
Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the Agentic AI systems scale or change context. These practices are designed to meet not only current standards like ISO/IEC 42001 for AI management systems, but also prepare for future regulatory evolutions. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.

### v3

@@ 7 Trust and Explainability in Agentic AI Systems
Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms. Figure 7 illustrates the layered relationship between trust mechanisms, explainability techniques, and evaluation facets. Trust-enabling mechanisms, such as transparency and human-in-the-loop (HIIL) oversight, provide the foundation for user confidence. These feed into explainability techniques like decision provenance graphs, which make agent behaviors interpretable. Finally, evaluation (Section 8) facets ensure that trust and explainability are quantifiable, enabling iterative improvements in system design.
@@ 7.1 Building Trust in Autonomous Agents
Transparency and Decision Traceability. Transparency is one of the core enablers of trust. For users to understand and evaluate AMAS decisions; the reasoning chains, decision states, and action triggers of agents must be made observable. Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent [128] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPT[129] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
@@
Behavioral Consistency and Bounded Autonomy. Trust requires predictability in models’ behaviour. Agents should follow defined roles, output in expected formats, and remain within delegated authority. For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow [130], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
Social Trust Cues and Language Behaviors. Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [141, 12, 142]. For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries . Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [143]. Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [134]. This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability. These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
@@ 7.2 Explainability Techniques in AMAS
The integration of trust-enabling mechanisms and explainability techniques is critical for the adoption of Agentic AI systems in high-stakes domains. While transparency, human oversight, and interpretable architectures improve user confidence, challenges in capturing system-level dynamics and scaling explainability remain. Future research should focus on developing standardized frameworks for multi-agent explainability, real-time interactive querying, and longitudinal trust evaluation to ensure robust and responsible deployment. A comparative summary of current techniques is shown in Table 6.
8 Evaluation of Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an AMAS performance and its real-world implications. Below, we discuss these metrics and summarize in Table 7.
Trustworthiness. This criteria evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to safety constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [153]. Some approaches combine such factors into an overall trust score. We can define a trustworthiness score $T$ for an Agentic AI system as :
@@ 8 Evaluation of Agentic AI Systems
Coordination. In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics measure how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
@@
9 Security and Privacy in Agentic AI Systems
@@ 9.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [180]. Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [181]. Among the foundational techniques employed are encryption , access control , adversarial defense , and runtime monitoring, each adapted to the unique demands of decentralized multi-agent environments.
@@
Encryption. Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [61]. Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption , and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
@@
Access control. Access conrol becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [53]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [182] and Attribute-Based Access Control (ABAC) [182] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
@@
Adversarial learning. Adversarial learning is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results. Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially designing outputs from one compromised agent propagating misleading information to others [183]. Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities. Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
@@
Runtime Monitoring. Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity . Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [184]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
@@ 9.2 Privacy-Preserving Techniques
The decentralized and interactive nature of AMAS introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory. To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
Differential privacy (DP). DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [185]. DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
@@
Privacy-by-design [187] principles are becoming central to the engineering of next-generation Agentic AI systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
10 Compliance to AI Governance and Policy in Agentic AI Systems
As Agentic AI systems grow more autonomous and capable, ensuring they remain compliant with regulations and governed by robust policies becomes imperative. Unlike traditional AI, these multi-agent or highly autonomous systems operate with persistent memory and complex decision-making loops, which introduces new risks around safety, ethics, and legality. Recent industry surveys indicate that enterprise leaders rank security, trust, compliance, and oversight as top concerns when deploying agentic AI at scale. To address these concerns, governance of agentic AI spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks. Below, we expand on each dimension and highlight emerging best practices for secure and compliant agentic AI deployments.
10.1 Regulatory standards
@@
Regulatory frameworks provide baseline requirements that all AI systems – including agent-based architectures – must meet. Key frameworks outline principles of transparency, accountability, safety, and fairness which are especially pertinent for autonomous multi-agent systems operating in high-stakes domains (e.g. healthcare, finance, transportation). Notable examples include:
  • NIST AI Risk Management Framework (AI RMF) [188], a voluntary U.S. framework emphasizing risk identification, mitigation, and trustworthy AI practices across the AI lifecycle. It calls for governance functions that ensure AI systems are transparent, accountable, and monitored for risks. Adopting the NIST AI RMF “Govern” and “Explain” functions helps organizations embed policies and documentation that align with trustworthiness goals.
  • EU AI Act [106], a comprehensive EU regulation (set to fully apply in 2026) that classifies AI systems by risk level. High-risk AI systems (a category that would likely include certain autonomous or agentic AI in critical fields) are subject to strict requirements: rigorous risk management, human oversight, transparency, and continuous monitoring. For example, the Act explicitly mandates ongoing risk assessment and traceability for high-risk AI, and it requires that such systems be designed with human-in-the-loop or human supervision capabilities (per Article 14 on oversight). Agentic AI developers must therefore build in features for traceability and human override to remain compliant with these upcoming rules.
  • Domain-Specific Regulations In addition to AI-focused standards, Agentic AI systems must comply with sectoral data and safety laws. For instance, an autonomous multi-agent healthcare assistant must adhere to health privacy laws like HIPAA [104], and any system handling personal data of EU citizens must follow GDPR [103] requirements for data protection by design. These regulations often require data minimization and user consent, which Agentic AI systems need to respect (e.g. by expiring or anonymizing stored user data to avoid retention of sensitive information).
10.2 Auditability
Auditability is critical for transparency and post-hoc accountability in agentic AI. Given the complexity of multi-agent decision-making (with potentially emergent or opaque interactions), organizations must be able to reconstruct and scrutinize the chain of events leading to any outcomes. Achieving robust auditability involves implementing the following practices:
  • Comprehensive Logging: Each decision, action, plan, and inter-agent message should be logged with rich metadata – including timestamps, the initiating agent’s identity/role, the context or prompt, and a rationale or justification for the action. Such detailed logs create a decision provenance trail, enabling auditors to trace how and why a particular outcome was reached.
  • Action Traceability: Techniques like decision provenance link each agent’s actions to preceding events or inputs, so one can follow the sequence of steps across agents. In complex agent workflows (e.g. a chain of a Planner agent delegating to a Coder agent, then to a Tester agent), traceability means every hand-off and decision node is recorded and can be reviewed in context. This helps identify where an error or policy violation might have occurred.
  • Role-Specific Audit Trails: Modern agentic frameworks often assign specialized roles to different agents (for example, in AutoGen or MetaGPT, agents may take on roles like Researcher, Coder, or Reviewer to collaborate on tasks). The audit system should capture actions at the granularity of each role. By tagging log entries with the agent role, one can detect anomalies or biases in specific functions – for instance, flagging if a “Reviewer” agent consistently approves outputs that the “Researcher” agent flagged as doubtful, or if a certain role’s actions lead to security issues
  • Immutable Logging: To ensure the integrity of audit logs (so that records cannot be tampered with after the fact), organizations are exploring append-only, tamper-proof logging mechanisms. For example, using blockchain or other immutable ledger technology to record multi-agent interactions can provide verifiable audit trails. This means any decision or message recorded in the system is cryptographically secured against alteration, which is valuable for compliance checks and forensic analysis in case of an incident.
Robust auditability not only helps with internal accountability but is also increasingly expected by regulators. Transparent record-keeping allows external auditors or regulators to verify compliance with regulations and ethical standards. In the event of a problematic outcome, detailed audit logs ensure that responsibility can be determined and corrective action taken. In fact, governance frameworks stress that organizations should implement such audit trails and explicable decision logs so that any harmful or biased result can be traced back and attributed appropriately. Overall, auditability in agentic AI builds the foundation for trust – both for internal stakeholders and for users or regulators – by making the system’s behavior observable and explainable after the fact.
10.3 Policy Enforcement
While regulations and audits provide oversight and after-the-fact analysis, policy enforcement mechanisms work in real-time to keep an Agentic AI system’s behavior within allowed bounds. Policy in this context refers to the codified rules and constraints that the AI agents must obey – covering operational limits, ethical guidelines, and security restrictions set by developers or governance boards. In practice, effective policy enforcement in agentic AI includes:
  • Orchestrator-Level Governance: Agentic AI systems typically have an orchestration layer or a meta-controller that coordinates multiple agents. This layer should be programmed to enforce global policies on what agents can do and under what conditions. For example, a policy might dictate that any plan an agent comes up with must receive approval from a supervisory agent (or a human) if it involves deleting or altering important data. The orchestration logic would then require that checkpoint before execution, overriding agent autonomy when necessary.
  • Memory and Data Retention Policies: To comply with privacy rules and avoid unchecked growth of an agent’s knowledge base, policies can enforce memory expiration or sanitization. Agents might be required to forget or redact certain information after a task is complete or after a time limit, preventing long-term retention of sensitive data. This guards against data misuse and helps comply with regulations like GDPR’s “data minimization” principle. For instance, if an agent system stores user queries or personal details, a policy could mandate purging those from the agent’s memory after fulfilling the query (unless explicitly allowed to retain).
  • Tool and Resource Access Control: Agentic AI often involves agents calling external tools, APIs, or databases to accomplish tasks. Role-Based Access Control (RBAC) and Attribute-Based Access Control (ABAC) are crucial for limiting what each agent (or agent role) is permitted to do. Under RBAC, each agent role has permissions aligned with its responsibilities – e.g. a “Coder” agent might be allowed to call a code compilation service but not to fetch customer data; a “Data Analyst” agent might read a database but not invoke external network calls. ABAC can add context-sensitive rules (attributes like time of day, the specific project, or the agent’s trust level) to further restrict actions.
  • Dynamic Sandboxing and Execution Constraints: Each agent’s actions (especially those that can change the environment or have side effects) should run in a sandbox or constrained context where possible. For example, if an agent is meant to write to a file, sandboxing can ensure it only writes to a designated directory and not to system-critical files. Similarly, agents might be prohibited from executing certain operations (like sending network requests to unapproved domains) through whitelists/blacklists coded into the environment. Dynamic sandboxing means these execution rules can adjust based on the agent’s role or current task, providing flexible yet controlled operation.
  • Real-time Monitoring and Intervention: Even with policies encoded, agents might encounter novel situations. Real-time monitoring systems serve as a safety net – they watch agent behavior and can halt or flag any activity that deviates from predefined ethical constraints or operational bounds. For instance, if an agent’s output is about to violate a known policy (say, it’s about to generate disallowed content or leak PII), a monitoring module can catch this and either censor the output or stop the agent’s process. Some approaches involve using formal verification and runtime checking tools (like TLA+ specifications or symbolic execution) to prove certain unsafe actions will not occur, or to continuously verify that invariants hold during execution.
10.4 Emerging Governance Practices
Because the field of Agentic AI is evolving rapidly, governance strategies are also adapting. Beyond the foundational controls above, organizations are adopting new best practices to strengthen oversight and ensure long-term compliance as these AI systems scale. Some of the emerging practices include:
  • AI Governance Boards: Many organizations are establishing dedicated governance boards or committees for AI oversight. These boards typically include stakeholders from legal, ethics, security, and domain experts. Their role is to continuously review the Agentic AI system’s objectives, assess risks, and update policies or strategies as needed. For example, a governance board might conduct regular audits of an AI’s decisions for bias or fairness, ensure alignment with the latest regulations, and approve any major changes to the system’s capabilities. This human layer of oversight creates accountability at the organizational level and sends a signal that the AI’s operations are being closely supervised by responsible parties.
  • Governance-as-Code: Inspired by the infrastructure-as-code paradigm in DevOps, governance-as-code means encoding governance policies and ethical rules directly into the software and configuration that drive the AI system. Instead of governance being only a manual or paperwork process, it becomes part of the system’s logic. For Agentic AI, this could involve using policy engines or rule-based systems that automatically check each agent action against a set of coded rules. For instance, a rule repository might specify “Agent of type X cannot use tool Y without permission Z” and this is enforced by the orchestration code.
  • Adaptive Governance Layers: Agentic AI systems may change over time – they might learn new skills, add more agents, or move into new deployment contexts. Adaptive governance refers to frameworks that evolve along with the AI system. This can include auto-updating rule sets or ML-driven policy adjustments. For example, if an agent starts operating in a new country, an adaptive governance layer might automatically incorporate that country’s AI regulations into the policy set. Or if a system update gives agents new capabilities, the governance rules adapt (or prompt human governors to adapt them) to cover those capabilities.
  • Standards Alignment and Certification: To have trust, organizations are aligning their Agentic AI governance with emerging international standards. A prime example is ISO/IEC 42001:2023 [189], and ISO/IEC 42005:2025 [101] the management system standard, which specify requirements for effective AI governance and risk management . ISO 42001 and 42005 emphasizes many of the same principles discussed above – such as transparency, accountability, bias mitigation, safety, and privacy in AI development and deployment. Forward-looking organizations are therefore proactively building those capabilities (e.g. implementing rigorous risk management pipelines, bias audits, etc.) so that by the time new laws fully kick in, their agentic AI will be in natural compliance.
By integrating these enforcement measures, Agentic AI systems can operate with much tighter alignment to intended guidelines and ethical norms. Policy enforcement translates high-level governance principles into day-to-day operational safeguards, reducing the chance of an autonomous agent straying outside its allowed domain or causing a compliance breach. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.